Television sets: What does it mean to refer to a 20 inch tv set or a 25 inch tv set? Such units refer to the diagonal of the screen. A 20 in set has a width of 16 inches what is the height?
What is the height of a 20 inch TV? _____inches

Answer by oscargut(686)

20^2=16^2+h^2
then
h^2=400-256=144
so height = 12 inches